# Constraining electromagnetic couplings of ultralight scalars from compact stars

### v1

@@ Abstract
If an ultralight scalar interacts with the electromagnetic fields of a compact rotating star, then a long-range scalar field is developed outside the star. The Coulomb-like profile of the scalar field is equivalent to an effective scalar charge on the star. In a binary star system, the scalar-induced charge would result in a long-range force between the stars, with the scalar field acting as the mediator. The scalar-photon interactions would modify Maxwell’s equations for electromagnetic fields in vacuum, resulting in a modified dispersion relation. This could be observed as an apparent redshift for photons emitted by such sources. The scalar field would also induce additional electric and magnetic fields and hence affect the electromagnetic energy radiated from such compact objects. A scalar field sourced by time-varying electromagnetic fields can also carry away energy from a compact star in the form of radiation, and hence contribute to its spin-down luminosity. We constrain the scalar-photon coupling from the measurements of the electromagnetic radiation of the compact star and its spin-down luminosity. We also project the prospective bounds on these couplings with future measurements of the apparent redshifts of compact stars and of the long-range force between two magnetars in a binary. We analyze the systems of the binary pulsar PSR J0737-3039, the Crab pulsar, the soft gamma repeater SGR 1806-20, and the gamma ray burst GRB 080905A. The bounds on the coupling can be significantly improved by future measurements of compact stars with large magnetic fields, experiments with better sensitivity, and precision clock measurements.
@@ I Introduction
Compact stars (NSs and magnetars) also serve as probes to search for the dark matter (DM) in the Universe Raffelt:1999tx ; Raffelt:2006cw ; Bramante:2023djs . Results from the Planck satellite suggest that the energy density of DM is about five times that of the visible matter Planck:2018nkj . The weakly-interacting massive particle (WIMP) motivated by the theory of supersymmetry has been one of the leading candidates for DM Jungman:1995df . However, constraints on WIMPs from direct detection experiments Roszkowski:2017nbc ; XENON:2018voc ; XENON:2020gfr ; LZ:2022lsv and the small scale structure of the Universe Bullock:2017xww motivate us to study alternative candidates for DM. Ultralight DM is one such promising candidate, where sub-eV mass range particles can account for the present relic density of the universe, at the same time staying consistent with the direct search experiments and cosmological observations Hu:2000ke ; Marsh:2015wka ; Hui:2016ltb ; Robles:2018fur ; Amin:2022pzv . If such a DM candidate has mass as low as $10^{-22}~{}\mathrm{eV}$, its de Broglie wavelength would be of the order of the size of a dwarf galaxy $(1-2~{}\mathrm{kpc})$. The number density of ultralight DM within this de Broglie wavelength is $10^{30}/\mathrm{cm^{3}}$ for the local DM density $\rho_{\odot}\sim 0.4~{}\mathrm{GeV}/\mathrm{cm^{3}}$. The presence of such a large number density implies that DM oscillates coherently in a wave-like manner or exhibits long-range behavior, potentially forming a Bose-Einstein condensate Boehmer:2007um ; Li:2013nal . The ultralight DM can be scalar Khmelnitsky:2013lxt ; Hui:2016ltb ; Badurina:2021lwr ; Tretiak:2022ndx , pseudoscalar Zhang:2016uiy ; Hook:2017psm ; Hook:2018iia ; Dror:2020zru ; Chadha-Day:2021szb ; Winch:2024mrt , vector Dror:2018pdh ; PPTA:2022eul ; Fedderke:2022ptm ; Chase:2023puj , or tensor Brito:2020lup ; Wu:2023dnp ; Guo:2023gfc ; some such particles are also motivated from string/M theory Witten:1984dg ; Battye:1994au ; Yamaguchi:1998gx ; Svrcek:2006yi ; Arvanitaki:2009fg .
@@
The phenomenology of ultralight scalar, pseudoscalar, vector, and tensor fields is remarkably diverse, and numerous studies over the years have explored their potential signatures in cosmic laboratories. The dilaton and axion field profiles in different string gravity models of BH have been discussed in Campbell:1990ai ; Campbell:1991kz ; Campbell:1991rz ; Mignemi:1992pm , where the field is sourced by the Chern-Simon and Gauss-Bonnet terms. In this paper, we consider the scenario where an ultralight scalar $\phi$ (which need not be the DM) interacts with the CP-even electromagnetic (EM) current $F_{\mu\nu}F^{\mu\nu}$. As we shall show later, such an interaction, in the presence of the large EM fields near the surface of a rotating compact star, leads to a long-range scalar field, [^3] Note that ultralight pseudoscalars such as axions may interact with the charge-parity (CP)-odd EM current $F_{\mu\nu}\tilde{F}^{\mu\nu}$. However, the resultant pseudoscalar field goes as $a\sim 1/r^{2}$ Harari:1992ea ; Mohanty:1993nh ; Krasnikov:1996bm ; McDonald:2019wou ; Sokolov:2022fvs ; Domcke:2023bat , i.e., it falls faster than the scalar field as one moves away from the source. The influence of an axion background on EM radiation is examined in McDonald:2019wou ; Sokolov:2022fvs ; Domcke:2023bat . $\phi\sim 1/r$. This scalar field, in turn, induces additional electric and magnetic fields around the source. We explore four kinds of effects of such a scalar field:
@@ II Long-range scalar field outside a compact star
$\mathbf{B}^{\mathrm{out}}_{(r>R)}=B_{0}R^{3}\Big{(}\frac{\cos\theta}{r^{3}}\hat{r}+\frac{\sin\theta}{2r^{3}}\hat{\theta}\Big{)}\,,$ (1)
@@
where $\phi$ denotes the scalar field and $g_{\phi\gamma\gamma}$ denotes the coupling constant of this scalar field with the EM fields of the star. The equation of motion of the scalar field can be obtained as
@@
This, in the Schwarzschild background, takes the form
$\frac{1}{r^{2}}\frac{\partial}{\partial r}\Big{[}(r^{2}-2Mr)\frac{\partial}{\partial r}\phi(r,\theta)\Big{]}+\frac{1}{r^{2}\sin\theta}\frac{\partial}{\partial\theta}\Big{[}\sin\theta\frac{\partial}{\partial\theta}\phi(r,\theta)\Big{]}=-g_{\phi\gamma\gamma}\frac{B^{2}_{0}R^{6}}{4r^{6}}(3\cos^{2}\theta+1)\,.$ (7)
Solving Eq. 7 by the Green’s function method, we obtain the scalar field profile as
@@
Thus, the rotating star has a long-range scalar “hair” associated with a charge $Q_{\phi}^{\mathrm{eff}}$. Though we have obtained Eqs. 8 and 9 for a massless scalar, our results would be valid as long as the Compton wavelength of the scalar is greater than the radius of the star, i.e., for $1/m_{\phi}\gtrsim R$, or $m_{\phi}\lesssim 1/R$.
Note that, for stars with a large angular velocity ($\Omega R\sim{\cal O}(1)$), the electric field outside the star cannot be neglected, and we need to solve the scalar field profile in the Kerr background, sourced by $\mathbf{B}^{2}-\mathbf{E}^{2}$ instead of only $\mathbf{B}^{2}$. A detailed analysis of this scenario is presented in Appendix A.
@@ III Scalar-induced EM fields from Maxwell’s equations
where the “$(0)$” corresponds to any quantity in the limit $g_{\phi\gamma\gamma}=0$. We keep the terms which are linear in $g_{\phi\gamma\gamma}$, and obtain $\partial_{\mu}F^{\mu\nu}_{\phi}=-g_{\phi\gamma\gamma}(\partial_{\mu}\phi)F^{\mu\nu}_{(0)}$ in the absence of source charge and current density of the plasma. This relation gives the expressions for the scalar induced electric $(\mathbf{E}_{\phi})$ and magnetic $(\mathbf{B}_{\phi})$ fields in terms of the background electric $(\mathbf{E}_{(0)})$ and magnetic $(\mathbf{B}_{(0)})$ fields as
@@
×$\nabla\times\mathbf{B}_{\phi}$ $=$ $\frac{\partial\mathbf{E}_{\phi}}{\partial t}-g_{\phi\gamma\gamma}\,\nabla\phi\times\mathbf{B}_{(0)}+g_{\phi\gamma\gamma}\left(\frac{\partial\phi}{\partial t}\right)\mathbf{E}_{(0)}\,,$ (11)
@@
×$\nabla\times\mathbf{E}_{\phi}$ $=$ $-\frac{\partial\mathbf{B}_{\phi}}{\partial t}\,.$ (12)
@@
$\begin{split}\Box\mathbf{B}_{\phi}=g_{\phi\gamma\gamma}\,(\nabla\phi\cdot\nabla)\,\mathbf{B}_{(0)}\,,\\ \Box\mathbf{E}_{\phi}=g_{\phi\gamma\gamma}\,(\nabla\phi\cdot\nabla)\,\mathbf{E}_{(0)}\,,\end{split}$ (13)
where we neglect terms which appear as two spatial derivatives of $\phi$ (since $\phi$ falls as $1/r$ and the derivatives of $\phi$ will fall even faster). In the limit $\Omega R\ll 1$, we solve Eq. 13 in the Schwarzschild background to obtain the scalar-induced magnetic field as [^4] The form of the scalar-induced magnetic field in the Kerr background is given in Appendix A.
$\begin{split}\mathbf{B}_{\phi}(r,\theta)\approx-\frac{g_{\phi\gamma\gamma}Q^{\mathrm{eff}}_{\phi}B_{0}R^{3}\cos\theta}{12M^{2}r^{2}}\,\hat{r}-\frac{g_{\phi\gamma\gamma}Q^{\mathrm{eff}}_{\phi}B_{0}R^{3}\pi}{64M^{3}r}\,\hat{\theta}\,,\end{split}$ (14)
@@
Note that the scalar-induced magnetic field $\mathbf{B}_{\phi}$ in Eq. 14 is actually proportional to $g^{2}_{\phi\gamma\gamma}$, as $Q^{\mathrm{eff}}_{\phi}$ itself is proportional to $g_{\phi\gamma\gamma}$. It falls as $1/r^{2}$ in the radial direction and $1/r$ in the angular direction, as compared to the background magnetic field $\mathbf{B}_{(0)}$ which falls as $1/r^{3}$. The scalar-induced electric field $\mathbf{E}_{\phi}$ can also be calculated in a similar manner. Since these scalar-induced EM fields scale quadratically with $g_{\phi\gamma\gamma}$, the deviations of the EM fields from their background values, as a result of their interactions with the scalar, are small.
@@ IV EM wave propagation in the background of a long-range scalar field
×$\nabla\times\mathbf{B}$ $=$ $\frac{\partial\mathbf{E}}{\partial t}-g_{\phi\gamma\gamma}\nabla\phi\times\mathbf{B}\,,$
@@
×$\nabla\times\mathbf{E}$ $=$ $-\frac{\partial\mathbf{B}}{\partial t}\,.$ (15)
@@
$\Box\mathbf{B}=g_{\phi\gamma\gamma}(\nabla\phi\cdot\nabla)\mathbf{B}\;.$ (16)
@@
$\omega^{2}-k^{2}=g_{\phi\gamma\gamma}k\nabla\phi\,,$ (17)
where $k=k(r)$. The group velocity of the photon becomes
$\frac{\partial\omega}{\partial k}=v_{g}=1+\frac{m^{2}_{\gamma}}{2\omega^{2}}+\mathcal{O}\Big{(}\frac{m^{3}_{\gamma}}{\omega^{3}}\Big{)}\,,$ (18)
where we define the scalar-induced photon mass as
$m_{\gamma}=|g_{\phi\gamma\gamma}\nabla\phi|/2\,.$ (19)
The correction in the photon group velocity appears with the fourth power of $g_{\phi\gamma\gamma}$, as $\nabla\phi\propto g_{\phi\gamma\gamma}$. Since $m^{2}_{\gamma}$ is a real positive quantity, The photon’s group velocity appears to be superluminal. However, this is not a problem since the group velocity of the photon in this case does not correspond to the speed of information propagation. Rather, it corresponds to the speed which appears as an analytical continuation of the pulse shape within the light cone Peters:1988mj ; Diener:1996mj ; Milonni:2005tos ; McDonald:2019wou . This is similar to the case of photon propagation in a dynamic axion background and in the absence of plasma, where the photon can exhibit a group velocity greater than one McDonald:2019wou .
From Eq. 17, we obtain the solution for $k$ as
$k=\omega\Big{(}1-\frac{m_{\gamma}}{\omega}+\frac{m^{2}_{\gamma}}{2\omega^{2}}\Big{)}.$ (20)
The contribution of this dispersion relation to the apparent redshift of the photon of wavelength $\lambda$, as measured in the asymptotically flat spacetime (i.e. at the observation point $r_{2}$) is then
$\delta z=\frac{\lambda(r_{2})-\lambda(r_{1})}{\lambda(r_{1})}\approx\frac{k(r_{1})-k(r_{2})}{k(r_{2})}\approx\frac{m_{\gamma}(r_{1})}{\omega}=\frac{g_{\phi\gamma\gamma}Q_{\phi}^{\mathrm{eff}}}{2\omega R^{2}},$ (21)
where $r_{1}$ represents the location close to the magnetar. We have assumed $m_{\gamma}(r_{2})\approx 0$ as the observer is far away from the magnetar (for example, at the Earth). From Eq. 21, the correction to the photon redshift would be significant if $m_{\gamma}(r_{1})$ is of the same order of magnitude as $\omega$. Moreover, the redshift as measured at different frequencies will be different, an indication of a non-trivial dispersion relation. Note that $\delta z$ is proportional to $g_{\phi\gamma\gamma}^{2}$, since $Q^{\mathrm{eff}}_{\phi}$ itself is proportional to $g_{\phi\gamma\gamma}$.

### v2

@@ Abstract
If an ultralight scalar interacts with the electromagnetic fields of a compact rotating star, then a long-range scalar field is developed outside the star. The Coulomb-like profile of the scalar field to the leading order is equivalent to an effective scalar charge on the star. In a binary star system, the scalar-induced charge would result in a long-range force between the stars, with the scalar field acting as the mediator. The scalar-photon interactions would modify Maxwell’s equations for electromagnetic fields in vacuum, resulting in a modified dispersion relation. This could be observed as an apparent redshift for photons emitted by such sources. The scalar field would also induce additional electric and magnetic fields and hence affect the electromagnetic energy radiated from such compact objects. A scalar field sourced by time-varying electromagnetic fields can also carry away energy from a compact star in the form of radiation, and hence contribute to its spin-down luminosity. We constrain the scalar-photon coupling from the measurements of the electromagnetic radiation of a compact star and from its spin-down luminosity, using the Crab pulsar, the soft gamma repeater SGR 1806-20, and the gamma ray burst GRB 080905A. We also project the prospective bounds on the coupling from future measurements of the long-range force between two compact stars in a binary such as PSR J0737-3039, and from the apparent redshifts of compact stars. Future advances in precision-clock sensitivity and targeted observations of stars with strong surface magnetic fields, large radii, and low-frequency emission can substantially tighten these coupling limits.
@@ I Introduction
Compact stars (NSs and magnetars) also serve as probes to search for the dark matter (DM) in the Universe Raffelt:1999tx ; Raffelt:2006cw ; Bramante:2023djs ; Caputo:2024oqc . Results from the Planck satellite suggest that the energy density of DM is about five times that of the visible matter Planck:2018nkj . The weakly-interacting massive particle (WIMP) motivated by the theory of supersymmetry has been one of the leading candidates for DM Jungman:1995df . However, constraints on WIMPs from direct detection experiments Roszkowski:2017nbc ; XENON:2018voc ; XENON:2020gfr ; LZ:2022lsv and the small scale structure of the Universe Bullock:2017xww motivate us to study alternative candidates for DM. Ultralight DM is one such promising candidate, where sub-eV mass range particles can account for the present DM relic density of the universe, at the same time staying consistent with the direct search experiments and cosmological observations Hu:2000ke ; Marsh:2015wka ; Hui:2016ltb ; Robles:2018fur ; Amin:2022pzv . If such a DM candidate has mass as low as $10^{-22}~\mathrm{eV}$, its de Broglie wavelength would be of the order of the size of a dwarf galaxy $(1-2~\mathrm{kpc})$. The number density of ultralight DM within this de Broglie wavelength is $10^{30}/\mathrm{cm^{3}}$ for the local DM density $\rho_{\odot}\sim 0.4~\mathrm{GeV}/\mathrm{cm^{3}}$. The presence of such a large number density implies that DM oscillates coherently in a wave-like manner or exhibits long-range behavior, potentially forming a Bose-Einstein condensate Boehmer:2007um ; Li:2013nal ; KumarPoddar:2019jxe ; DiGiovanni:2022mkn ; Davoudiasl:2024grq ; Poddar:2024thb ; Dutta:2024vzw . The ultralight DM can be scalar Khmelnitsky:2013lxt ; Hui:2016ltb ; Badurina:2021lwr ; Tretiak:2022ndx ; Matos:2023usa , pseudoscalar Zhang:2016uiy ; Hook:2017psm ; Hook:2018iia ; Dror:2020zru ; Chadha-Day:2021szb ; Winch:2024mrt ; Li:2024bbe , vector Dror:2018pdh ; PPTA:2022eul ; Fedderke:2022ptm ; Chase:2023puj ; Chowdhury:2023xvy , or tensor Brito:2020lup ; Wu:2023dnp ; Guo:2023gfc ; some such particles are also motivated from string/M theory Witten:1984dg ; Battye:1994au ; Yamaguchi:1998gx ; Svrcek:2006yi ; Arvanitaki:2009fg .
@@
The phenomenology of ultralight scalar, pseudoscalar, vector, and tensor fields is remarkably diverse, and numerous studies over the years have explored their potential signatures in cosmic laboratories. The dilaton and axion field profiles in different string gravity models of BH have been discussed in Campbell:1990ai ; Campbell:1991kz ; Campbell:1991rz ; Mignemi:1992pm , where the field is sourced by the Chern-Simon and Gauss-Bonnet terms. In this paper, we consider the scenario where an ultralight scalar $\phi$ (which need not be the DM) interacts with the CP-even electromagnetic (EM) current $F_{\mu\nu}F^{\mu\nu}$. As we shall show later, such an interaction, in the presence of the large EM fields near the surface of a rotating compact star, leads to a long-range scalar field, [^3] Note that ultralight pseudoscalars such as axions may interact with the charge-parity (CP)-odd EM current $F_{\mu\nu}\tilde{F}^{\mu\nu}$. However, the resultant pseudoscalar field goes as $a\sim\cos\theta/r^{2}$ Mohanty:1993nh , where $\theta$ is the polar angle for a rotating magnetized compact star, i.e., it falls faster than the scalar field as one moves away from the source. The influence of an axion background on EM radiation is examined in Harari:1992ea ; Krasnikov:1996bm ; McDonald:2019wou ; Sokolov:2022fvs ; Domcke:2023bat . $\phi\sim 1/r$. This scalar field, in turn, induces additional electric and magnetic fields around the source. We explore four kinds of effects of such a scalar field:
@@
In this paper, we do not explore the detailed mechanisms of scalar mass generation, since our focus is not on model building or ultraviolet completion. However, it is known that such ultralight scalars can naturally acquire mass in several theoretical frameworks. These include clockwork mechanism Giudice:2016yja ; Wood:2023lis , Planck-suppressed operators from string theory or quantum gravity Hui:2016ltb ; Hubisz:2024hyz ; Banerjee:2022wzk , and non-perturbative effects such as instantons Kitano:2021fdl . Ultralight scalars also commonly appear in extra-dimensional compactifications Anchordoqui:2023tln , or scale–invariant symmetry breaking scenarios Ferreira:2020fam .
@@ II Long-range scalar field outside a compact star
[…]
@@
where $\phi$ denotes the scalar field and $g_{\phi\gamma\gamma}$ denotes the effective coupling of this scalar field with the EM fields of the star. Note that $g_{\phi\gamma\gamma}$ can have either sign, as determined by the nature of the charged particles running in the triangle loop diagrams responsible for the effective scalar-photon interaction. When the loop contribution is dominated by fermions such as leptons or heavy quarks, the coupling $g_{\phi\gamma\gamma}$ is positive. However, if the loop is dominated by a $W$ boson, the coupling would acquire a negative sign Flambaum:2024zyt .
The equation of motion of the scalar field can be obtained as
@@
Solving Eq. 6 by the Green’s function method in the Schwarzschild background, we obtain the scalar field profile as
@@
Here, $M$ is the mass of the compact star. Note that the scalar-induced effective charge, $Q^{\mathrm{eff}}_{\phi}$, can be either positive or negative, depending on the sign of the coupling $g_{\phi\gamma\gamma}$. The dependence of $Q^{\text{eff}}_{\phi}$ on $M$ in Eq. 8 is a result of general relativistic corrections to the scalar field profile, arising from the Schwarzschild metric. In flat spacetime, the effective charge simplifies to $Q^{\text{eff}}_{\phi}\sim g_{\phi\gamma\gamma}B_{0}^{2}R^{3}$. This form arises from the term $F_{\mu\nu}F^{\mu\nu}\propto|\mathbf{B}|^{2}$ in the scalar source charge density $\rho_{\phi}$, under the assumptions of slow rotation and a magnetically dominated regime $|\mathbf{B}|\gg|\mathbf{E}|$.
The $1/M^{3}$ dependence of $Q_{\phi}^{\mathrm{eff}}$ in Eq. 8 can be motivated as follows. The scalar field profile is obtained by solving Eq. 6, where the $M$- dependence is implicit through the d’Alembertian operator. This operator gives general-relativistic $(M/R)$-corrections to the flat spacetime solution. Given that the flat spacetime solution is $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{3}$ while Eq. 6 forces us to have $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{6}$, dimensional analysis tells us that the $(M/R)$-dependence should be of the form $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{3}(R/M)^{3}$. Explicit calculation also gives the same dependence on $M$.
Note that while deriving the scalar field profile above, the metric inside the star was taken to be Schwarzschild for analytical simplicity. This approximation is justified since replacing the Schwarzschild interior with a full TOV (Tolman-Oppenheimer-Volkoff) solution would modify the scalar charge by only a few percent for a typical NS. One gets $\Delta Q_{\phi}/Q_{\phi}\sim\mathcal{O}(C\,p/\rho)$, where $C\equiv M/R$ denotes the stellar compactness, and $p$ and $\rho$ represent the pressure and density inside the star, respectively. Such a small correction ($\Delta Q_{\phi}/Q_{\phi}\lesssim 10\%$, for typical NS parameters) is well within the astrophysical uncertainties in the stellar radius and magnetic field, making the approximation adequate for the level of accuracy considered in this work.
Equations 7 and 8 indicate that the rotating star has a long-range scalar “hair” associated with a charge $Q_{\phi}^{\mathrm{eff}}$. Though we have obtained these results for a massless scalar, our results would be valid as long as the Compton wavelength of the scalar is greater than the radius of the star, i.e., for $1/m_{\phi}\gtrsim R$, or $m_{\phi}\lesssim 1/R$.
We use GRB 080905A as a benchmark and write Eq. 8 as
$Q_{\phi}^{\mathrm{eff}}=2.4\times 10^{38}\Big(\frac{g_{\phi\gamma\gamma}}{10^{-15}~\mathrm{GeV^{-1}}}\Big)\Big(\frac{B_{0}}{3.93\times 10^{16}~\mathrm{G}}\Big)^{2}\Big(\frac{R}{10~\mathrm{km}}\Big)^{6}\Big(\frac{1.4~M_{\odot}}{M}\Big)^{3}.$ (9)
Note that, for stars with a large angular velocity ($\Omega R\sim{\cal O}(1)$), the electric field outside the star cannot be neglected, and we need to solve the scalar field profile sourced by $\mathbf{B}^{2}-\mathbf{E}^{2}$ instead of only $\mathbf{B}^{2}$. A detailed analysis of this scenario is presented in Appendix A.
@@ III Scalar-induced EM fields from Maxwell’s equations
where the “(0)(0)” corresponds to any quantity in the limit $g_{\phi\gamma\gamma}=0$. We keep the terms which are linear in $g_{\phi\gamma\gamma}$, and obtain $\partial_{\mu}F^{\mu\nu}_{\phi}=-g_{\phi\gamma\gamma}(\partial_{\mu}\phi)F^{\mu\nu}_{(0)}$ in the absence of source charge and current density of the plasma. This relation gives the expressions for the scalar induced electric $(\mathbf{E}_{\phi})$ and magnetic $(\mathbf{B}_{\phi})$ fields in terms of the background electric $(\mathbf{E}_{(0)})$ and magnetic $(\mathbf{B}_{(0)})$ fields as
@@
$\nabla\times\mathbf{B}_{\phi}$ $=$ $\frac{\partial\mathbf{E}_{\phi}}{\partial t}-g_{\phi\gamma\gamma}\,\nabla\phi\times\mathbf{B}_{(0)}+g_{\phi\gamma\gamma}\left(\frac{\partial\phi}{\partial t}\right)\mathbf{E}_{(0)}\,$, (11)
@@
$\nabla\times\mathbf{E}_{\phi}$ $=$ $-\frac{\partial\mathbf{B}_{\phi}}{\partial t}\,.$ (12)
@@
[…]
where we neglect terms which appear as two spatial derivatives of $\phi$ (since $\phi$ falls as $1/r$ and the derivatives of $\phi$ will fall even faster). In the limit $\Omega R\ll 1$, we solve Eq. 13 in the Schwarzschild background to obtain the scalar-induced magnetic field as [^4] The form of the scalar-induced magnetic field for large angular velocity is given in Appendix A.
[…]
@@
Note that the scalar-induced magnetic field $\mathbf{B}_{\phi}$ in Eq. 14 is actually proportional to $g^{2}_{\phi\gamma\gamma}$, as $Q^{\mathrm{eff}}_{\phi}$ itself is proportional to $g_{\phi\gamma\gamma}$. It falls as $1/r^{2}$ in the radial direction and $1/r$ in the angular direction, as compared to the background magnetic field $\mathbf{B}_{(0)}$ which falls as $1/r^{3}$ both in radial and angular directions. The scalar-induced electric field $\mathbf{E}_{\phi}$ can also be calculated in a similar manner. Since these scalar-induced EM fields scale quadratically with $g_{\phi\gamma\gamma}$, the deviations of the EM fields from their background values, as a result of their interactions with the scalar, are small.
@@ IV EM wave propagation in the background of a long-range scalar field
$\nabla\times\mathbf{B}$ $=$ $\frac{\partial\mathbf{E}}{\partial t}-g_{\phi\gamma\gamma}\nabla\phi\times\mathbf{B}\,$,
@@
$\nabla\times\mathbf{E}$ $=$ $-\frac{\partial\mathbf{B}}{\partial t}\,.$ (15)
@@
$\begin{split}\Box\mathbf{B}=g_{\phi\gamma\gamma}(\nabla\phi\cdot\nabla)\mathbf{B}\,,\\ \Box\mathbf{E}=g_{\phi\gamma\gamma}(\nabla\phi\cdot\nabla)\mathbf{E}\,.\end{split}$ (16)
@@
$\omega^{2}=k^{2}-ig_{\phi\gamma\gamma}(\nabla\phi\cdot\mathbf{k})\,$, (17)
where the imaginary component implies that the wavenumber of photon must have a complex value in the presence of a scalar background.
We obtain the expression for the group velocity from Eq. 17 as
$v_{g}=\frac{d\omega}{dk}=\frac{2k-im_{\gamma}}{2\omega(k)},$ (18)
where we define the scalar-induced photon mass $m_{\gamma}=|g_{\phi\gamma\gamma}\nabla\phi|$, assuming the photon propagation is radial. We solve Eq. 17 for $k$ as
$k=k_{R}+ik_{I}=\frac{\sqrt{4\omega^{2}-m_{\gamma}^{2}}}{2}+\frac{im_{\gamma}}{2},$ (19)
where $k_{R}=\sqrt{4\omega^{2}-m_{\gamma}^{2}}/2$ and $k_{I}=m_{\gamma}/2$. Note that $k_{R}$ is a real quantity provided $\omega>m_{\gamma}/2$. Using Eq. 19, we obtain the expression for the group velocity in Eq. 18 as
$v_{g}=\Big(1-\frac{m_{\gamma}^{2}}{4\omega^{2}}\Big)^{\frac{1}{2}}.$ (20)
In the region $\omega\gtrsim m_{\gamma}/2$, the group velocity is subluminal and can be approximated as $v_{g}\approx 1-(m_{\gamma}^{2}/8\omega^{2})$. We observe that the group velocity of the photon remains unchanged at order $\mathcal{O}(g^{2}_{\phi\gamma\gamma})$ since $\nabla\phi\propto g_{\phi\gamma\gamma}$; it receives corrections only at $\mathcal{O}(g^{4}_{\phi\gamma\gamma})$. If there is no scalar-photon coupling, then $m_{\gamma}\rightarrow 0$ and $v_{g}$ becomes unity.
The wavelength of the propagating wave is governed by the real part of the wavenumber, $k_{R}$, in the region $\omega>m_{\gamma}/2$. In this region, $k_{R}\simeq\omega-(m^{2}_{\gamma}/8\omega)$. The contribution of this dispersion relation to the apparent redshift of the photon of wavelength $\lambda$, as measured in the asymptotically flat spacetime (i.e., at the observer point $r_{2}$), is then
$\delta z=\frac{\lambda(r_{2})-\lambda(r_{1})}{\lambda(r_{1})}\approx\frac{k_{R}(r_{1})-k_{R}(r_{2})}{k_{R}(r_{2})}\approx\frac{m^{2}_{\gamma}}{8\omega^{2}}\approx\frac{g^{4}_{\phi\gamma\gamma}B^{4}_{0}R^{8}}{48^{2}\times 8M^{6}\omega^{2}},$ (21)
where, in the last approximations, we consider the relevant magnitude of the apparent redshift and assume that the wave propagates along the direction of $\nabla\phi$, i.e., radially. Here $r_{1}$ represents a location close to the magnetar. We have assumed $m_{\gamma}(r_{2})\approx 0$ as the observer is far away from the magnetar (for example, at the Earth). From Eq. 21, the correction to the photon redshift would be significant if $m_{\gamma}(r_{1})$ is of the same order of magnitude as $\omega$. Moreover, the redshift as measured at different frequencies will be different, an indication of a non-trivial dispersion relation.